OpenStudy (anonymous):

derivative of sin(x+3)

OpenStudy (anonymous):

cos(x+3) * d/dx (x+3) yields cos(x+3)

OpenStudy (anonymous):

craneology is right

OpenStudy (anonymous):

if i need to rewrite it as a sum identity for sine, would it be.. sin(x+3) = sinx*cos3 + cosx*sin3 ?

OpenStudy (anonymous):

yes, you are right! sin(x+3) = sinx*cos3 + cosx*sin3 =)